# Improving Code Performance Using LLMs in Zero-Shot: RAPGen

### v2

@@ I Introduction
Performance bugs are inefficient code snippets in software that can unnecessarily waste time and resources. Unlike functional bugs, performance bugs do not usually cause system failure. They tend to be harder to detect [Attariyan et al.(2012), Dean et al.(2014a), Dean et al.(2014b), Jovic et al.(2011)] and fix compared to functional bugs [Nistor et al.(2013), Song and Lu(2014)] and are usually fixed by expert developers [Chen et al.(2019)] and often manifest through large inputs or specific execution configurations [Nistor et al.(2013), Olivo et al.(2015)]. Therefore, they can sometimes go undetected for a long periods of time [Dean et al.(2014b), Jovic et al.(2011)].  As a result, better tool support is needed to fix performance bugs, specially for novice developers. For such a tool to be applicable in practice, it should also support a wide-range of performance bugs. However, the existing fixing approaches usually target specific kinds of performance bugs, such as repeated computations [Della Toffola et al.(2015)], software misconfigurations [Iqbal et al.(2021)], loop inefficiencies [Nistor et al.(2013)]. The majority of these approaches are also rule-based analyzers, with high maintenance cost [Bielik et al.(2017)]. Recently, to address this challenge, Garg et al. developed a deep-learning based performance fix tool called DeepDev-PERF [Garg et al.(2022)] that supports a variety of performance bugs and does not require maintaining explicit rules or an expert system. Our approach builds on that work by showing the cost of building and maintaining such a tool can be further improved through the use of prompt engineering as opposed to intensive fine-tuning.
@@
With the rising adoption of tools such as Github copilot [Chen et al.(2021)] by developers, large language models are showing a promising success as productivity tools. Additionally, a promising paradigm for few-shot and zero-shot learning has emerged where a pre-trained large language model is adapted to various downstream tasks by conditioning it on natural language prompts. This paradigm has also been applied to a variety of tasks in Software Engineering, including the task of bug-fixing [Pearce et al.(2021b), Xia and Zhang(2022)]. Building upon this prior work, we present an approach called RAPGen that uses large language models (LLMs) in zero-shot, for the task of generating fixes for performance bugs using prompt engineering. Given a line of code that contains a performance bug (i.e. the buggy line), we compare this line with a pre-constructed knowledge-base to retrieve a prompt instruction that can be used to convey what change needs to be made to a LLM. The index contains patterns of buggy API usages extracted from a data-set of past performance fixes, along with an instruction explaining the fix transformation. Given a method with a performance issue and the corresponding buggy line within the method, we first retrieve an instruction from the knowledge-base and then construct an input prompt for the model, containing the buggy method, followed by the instructive comment and the method signature, coaxing the model to output the fixed version of the method. This type of zero-shot bug-fixing experience is desirable because (1) it avoids an expensive fine-tuning step, (2) alleviates the need to find a large high-quality labeled data-set of performance fixes for training such a model, and (3) fixes a wide-range of performance problems and could easily be extended to more bugs (by growing the number of commits used to build the knowledge-base) or other languages (by collecting relevant commits from other languages), without needing to train a separate model.
Evaluation and Effectiveness. We evaluated our approach on the most recent performance bug fix data set used in the DeepDev-PERF study [Garg et al.(2022)]. Experimental results show that our approach is indeed able to generate correct suggestions for 60% of the dataset, often within the very first suggestion, and out-performs DeepDev-PERF and other baselines for single method performance fixes. While designing an optimal prompt is still an open problem and remains elusive as ever, we also compare our approach with various prompt variations to demonstrate our design choices and our experiments with recent ideas such as reasoning-based prompting [Wei et al.(2022)]. Furthermore, through our "In-the-wild" evaluation, we show real-world evidence that RAPGen generates suggestions that lead to actual performance improvements to real .NET service codebases within a large company that are currently in production. These changes were validated and adopted by real developers.
@@ II Motivating Example
In this section, we provide an overview of RAPGen using a motivating example. Figure 1 shows a C# code method with a performance bug in the form of an expensive LINQ [Pialorsi and Russo(2007)] query (highlighted in red). The query iterates over a collection, finding all entries that match the predicate passed to the LINQ method Where, finally using FirstOrDefault to get the first such entry for further processing. While other languages have equivalent concepts, LINQ is specific to C# and is known to inherent allocations associated with its use. As a result, LINQ usage on the application’s hot-path can lead to excessive allocations, which can cause unwanted symptoms such as spikes in garbage collection (GC), depriving the application of CPU resources, and reducing its throughput.
Figure 4 shows the steps in our approach to generate fixes for a given buggy method and expensive line of code. To fix this code, we first compare the line of code containing the bug (that can be found easily using a profiler [Gayhardt(2019)]), by abstracting out project-specific details and querying a pre-built knowledge-base containing past buggy API usages, which includes the buggy usages of FirstOrDefault. This gives us a prompt instruction in English language telling how this bug should be fixed, which we then use to construct the prompt as shown in Figure 2. The prompt includes the buggy method itself as a comment, followed by the prompt instruction, telling the model to re-write the preceding method without the FirstOrDefault call, implying that it should be replaced with cheaper alternatives, which in this case would be an inlined foreach loop. To further ensure that the the model completion contains the commented method, the prompt also includes the signature of the buggy method, followed by an open curly brace, at the end of prompt. This is to coax the model into outputting a completion of that same method. Figure II shows the completion generated by the LLM to this prompt, which indeed replaces the LINQ query with an explicit foreach loop.
@@ III-A Code Transformation Knowledge-base
Data Collection We first collect all repositories with $\geq$5 stars on GitHub, whose primary language is C# and had a commit within the last 5 years. We crawl the commit history of the main branch in each of these repos. A commit typically contains a commit message and a changeset or diff representing the difference between current and previous version of files modified within that commit. From these repos, we collect the set of commits whose titles begin with "PERF:" or "[PERF]". This is a convention [com(2019)] that’s often used by developers to indicate that their commit contains primarily performance-related changes. This results in $\sim$1500 commits with clean code changes pertaining to performance.
@@ III-B Generating the Prompt
We are given buggy method, $m_{b}$, containing an expensive line of code, $l$, which causes the performance bug. We first abstract the expensive line using Algorithm 1 and get the abstracted pattern $p_{b}$. We then perform a lookup in our knowledge-base using the $p_{b}$, as the query. This results in a list of $K_{p}$ code change and instruction triples, i.e. $[(m_{k},m^{\prime}_{k},t_{k})]_{k=1}^{K_{p}}$, where $m_{k}$ and $m^{\prime}_{k}$ are the before and after versions of a method, where a similar buggy line was fixed and $t_{k}$ is the code transformation instruction constructed based on changes made to the identifiers in the before-after pair. Since we want to find the entry that’s closest to $m_{b}$, we use a code-search technique  [Luan et al.(2019)] to rank each before method, $m_{i}$, based on similarity to $m_{b}$. We take the entry corresponding to the most similar before method, $(m_{j},m^{\prime}_{j},t_{j})$ and construct the prompt using the retrieved code transformation instruction $t_{j}$.
@@ III-C Generating Fixes
Generating the fix itself is relatively simple. We use the constructed prompts to generate completions using the gpt-3.5-turbo [ope(2023)] model through OpenAI’s REST API. Since it is difficult to tell the model when to stop, it will likely go on outputting code even after it has generated the fixed method. We parse out the completion corresponding to the method by extracting the code until curly braces are balanced and discard the rest of the output.
@@ IV-A Experimental Setup
Dataset We leverage the dataset containing past performance fixes from the DeepDev-PERF study [Garg et al.(2022)]. This dataset consists of 132 instances of performance fixes, which were manually reviewed and confirmed to be performance-related changes by performance experts after examining a broad set of C# commits on GitHub. Furthermore, it was shown that this dataset covers a wide-range of performance bugs. Therefore, a good accuracy over this dataset would demonstrate the effectiveness of a model for this task.Each example in this dataset is of the form $(m,m^{\prime})$, where $m$ and $m^{\prime}$ are the before and after versions, respectively. These are also mostly single method performance improvement fixes, making them ideal for our scenario.
Adding Bug Localization Information Our approach assumes that bug localization information, i.e. the expensive line, is included as part of the input. In practice, one can leverage performance bug localization using profiling data to find the expensive line [Garg et al.(2021)]. However, for the purpose of evaluating our approach, we used a heuristic to localize bugs in the above dataset. We examine the diff between the before and after methods, $m$ and $m^{\prime}$ and take the first line $l_{b}\in m$ that does not match the corresponding line in $m^{\prime}$, when compared in order, as the buggy line.
@@
  • CodeBLEU: Our final metric is CodeBLEU [Ren et al.(2020)], which is a variation of the popular NLP metric, BLEU [Papineni et al.(2002)]. In addition to n-gram matching, it also takes code characteristics like abstract syntax tree (AST) and data-flow similarity into account, when comparing programs. Similar to the DeepDev-PERF study, we use the hyper-parameters that had the highest correlation to human scores in the CodeBLEU study, i.e. $\alpha,\beta,\gamma,\delta=0.1,0.1,0.4,0.4$.
@@
  • Closest Match Top-K Accuracy % Due to there being too many suggestions to go through one by one, we use the state of the art code search technique, Aroma [Luan et al.(2019)], to find the suggestion closest to the ground truth. The most similar suggestion is shown to two performance experts, who are then asked whether they consider the model suggestion to be equivalent or better that the developer fix in terms of performance. We consider a suggestion to be correct, if it is found to be correct by both experts. This allows us to get a lower-bound on the model’s actual Top-K accuracy, if one were to exhaustively go through every suggestion.
@@
  • Reasoning-based Prompt Message. Recent work on prompt engineering suggests letting LLMs generate an intermediate reasoning about a problem can greatly improve performance on a given task [Wei et al.(2022)]. As our final variant, we designed a reasoning-based prompt that follows the structure in Figure 5(c). This prompt also includes the buggy method, $m_{b}$, followed by an incomplete comment telling the model what call is being expensive, and coaxing the model to reason about the problem and then generate a possible fix. The main idea is that we first get the model to reason about the problem and come up with a description of the possible fix by completing the prompt message itself, before it generates the fixed code. We do this by giving the model the buggy method, followed by the prompt message. We let the model complete the rest of the comment (i.e. continuing in place of the ellipsis) and then have it generate a fix as well. The prompt follows structure in Figure 5(c).
@@ IV-B Comparing with Other Prompt Variants
Figure 10: An email sent as part of our "In-the-wild" Evaluation. The email contains a fix for a performance bottleneck found using Code Optimizations [cod(2023)] in a Microsoft service. Along with the fix, we also include information regarding the bottleneck, i.e. the bottleneck method and the calling parent method (both highlighted in yellow) and current CPU usage of the bottleneck. At the end of the email, we ask the users to answer a few questions regarding the fix (Section V-A2).
Effect of extracting instruction from retrieved example. Prior work has shown that providing an example in a few shot setting usually results in better performance in a downstream task than providing an instruction[Brown et al.(2020)]. It likely has less to do with the example but more with specificity of the instructions The Retrieval-based One-Shot prompt gives the model the same retrieved before-after pair that the instruction is derived from. However, comparing the results of RAPGen with the One-Shot prompt, we can see that going the extra step of extracting an instruction from the retrieved before-after pair improves the results significantly. We believe that this may be because, given the raw before-after method pair, the model has to perform the diff implicitly and infer the instruction by itself, before applying the change to the unseen method. It is possible for it misunderstand the change and make mistakes in understanding the nature of the fix. We see that our prompts find more exact matches (by $>$10%) and achieves a higher Top-K accuracy score (by $>$20%), compared to retrieval prompts. For the example shown in Figure 9, both the One Shot and RAPGen prompts were able to generate the right fix in our evaluation. However, the One Shot prompt took almost twice as many attempts before arriving at the right fix, compared to RAPGen. The likely reason RAPGen is able to generate a correct suggestion more quickly may be due to a possible fix being readily available as part of the instruction, without the model having to perform in-context learning from a raw code change example and infer the underlying transform.
@@ IV-C Comparison with State-of-the-art
Finally, we compare our approach with the current state of the art approach for performance bug fixing for C#: DeepDev-PERF  [Garg et al.(2022)]. DeepDev-PERF uses a sequence to sequence transformer model, BART [Lewis et al.(2020)] and finetunes it for the task of generating performance bugs. The model can output entire patches, as opposed to just completions, allowing it to generate fixes spanning the entire class. In contrast, our approach is intended to fix single method bugs only. Therefore, we only compare our model with DeepDev-PERF’s ability to fix single method fixes, using the dataset they provide.
@@ V-A1 Finding real world problems to fix
We start with 50 codebases, internal to a large company, that are associated with existing services running in the cloud. Each of these services is being profiled by Microsoft Application Insights Profiler [Gayhardt(2019)]. The collected profiles are then analyzed by a performance bug detection service called Code Optimizations [cod(2023)] that continuously analyzes profiler traces from a given application in the cloud and finds performance bottlenecks.
Each bottleneck is of the form $(b,p)$, where $b$ in the bottleneck method that’s usually an API method and $p$ the calling method from the user’s codebase. We look at the insights generated by Code Optimizations over CPU profiler traces. For each service, we find bottlenecks $(b,p)$ that meet the following criteria:

### v4

# RAPGen: An Approach for Fixing Code Inefficiencies in Zero-Shot

@@ Abstract
Index Terms:
Software Performance, Bug Repair, Large Language Models, AI for SE.
@@ I Introduction
Performance bugs are inefficient code snippets in software that can unnecessarily waste time and resources. Unlike functional bugs, performance bugs do not usually cause system failure. They tend to be harder to detect [1, 2, 3, 4] and fix compared to functional bugs [5, 6] and are usually fixed by expert developers [7] and often manifest through large inputs or specific execution configurations [5, 8]. Therefore, they can sometimes go undetected for a long periods of time [3, 4].  As a result, better tool support is needed to fix performance bugs, specially for novice developers. For such a tool to be applicable in practice, it should also support a wide-range of performance bugs. However, the existing fixing approaches usually target specific kinds of performance bugs, such as repeated computations [9], software misconfigurations [10], loop inefficiencies [5]. The majority of these approaches are also rule-based analyzers, with high maintenance cost [11]. Recently, to address this challenge, Garg et al. developed a deep-learning based performance fix tool called DeepDev-PERF [12] that supports a variety of performance bugs and does not require maintaining explicit rules or an expert system. Our approach builds on that work by showing the cost of building and maintaining such a tool can be further improved through the use of prompt engineering as opposed to intensive fine-tuning.
@@
With the rising adoption of tools such as Github copilot [13] by developers, large language models are showing a promising success as productivity tools. Additionally, a promising paradigm for few-shot and zero-shot learning has emerged where a pre-trained large language model is adapted to various downstream tasks by conditioning it on natural language prompts. This paradigm has also been applied to a variety of tasks in Software Engineering, including the task of bug-fixing [14, 15]. Building upon this prior work, we present an approach called RAPGen that uses large language models (LLMs) in zero-shot, for the task of generating fixes for performance bugs using prompt engineering. Given a line of code that contains a performance bug (i.e. the buggy line), we compare this line with a pre-constructed knowledge-base to retrieve a prompt instruction that can be used to convey what change needs to be made to a LLM. The index contains patterns of buggy API usages extracted from a data-set of past performance fixes, along with an instruction explaining the fix transformation. Given a method with a performance issue and the corresponding buggy line within the method, we first retrieve an instruction from the knowledge-base and then construct an input prompt for the model, containing the buggy method, followed by the instructive comment and the method signature, coaxing the model to output the fixed version of the method. This type of zero-shot bug-fixing experience is desirable because (1) it avoids an expensive fine-tuning step, (2) alleviates the need to find a large high-quality labeled data-set of performance fixes for training such a model, and (3) fixes a wide-range of performance problems and could easily be extended to more bugs (by growing the number of commits used to build the knowledge-base) or other languages (by collecting relevant commits from other languages), without needing to train a separate model.
Evaluation and Effectiveness. We evaluated our approach on the most recent performance bug fix data set used in the DeepDev-PERF study [12]. Experimental results show that our approach is indeed able to generate correct suggestions for 60% of the dataset, often within the very first suggestion, and out-performs DeepDev-PERF and other baselines for single method performance fixes. While designing an optimal prompt is still an open problem and remains elusive as ever, we also compare our approach with various prompt variations to demonstrate our design choices and our experiments with recent ideas such as reasoning-based prompting [16]. Furthermore, through our "In-the-wild" evaluation, we show real-world evidence that RAPGen generates suggestions that lead to actual performance improvements to real .NET service codebases within a large software company that are currently in production. These changes were validated and adopted by real developers.
@@ II Motivating Example
In this section, we provide an overview of RAPGen using a motivating example. Figure 1 shows a C# code method with a performance bug in the form of an expensive LINQ [17] query (highlighted in red). The query iterates over a collection, finding all entries that match the predicate passed to the LINQ method Where, finally using FirstOrDefault to get the first such entry for further processing. While other languages have equivalent concepts, LINQ is specific to C# and is known to inherent allocations associated with its use. As a result, LINQ usage on the application’s hot-path can lead to excessive allocations, which can cause unwanted symptoms such as spikes in garbage collection (GC), depriving the application of CPU resources, and reducing its throughput.
Figure 4 shows the steps in our approach to generate fixes for a given buggy method and expensive line of code. To fix this code, we first compare the line of code containing the bug (that can be found easily using a profiler [18]), by abstracting out project-specific details and querying a pre-built knowledge-base containing past buggy API usages, which includes the buggy usages of FirstOrDefault. This gives us a prompt instruction in English language telling how this bug should be fixed, which we then use to construct the prompt as shown in Figure 2. The prompt includes the buggy method itself as a comment, followed by the prompt instruction, telling the model to re-write the preceding method without the FirstOrDefault call, implying that it should be replaced with cheaper alternatives, which in this case would be an inlined foreach loop. To further ensure that the the model completion contains the commented method, the prompt also includes the signature of the buggy method, followed by an open curly brace, at the end of prompt. This is to coax the model into outputting a completion of that same method. Figure II shows the completion generated by the LLM to this prompt, which indeed replaces the LINQ query with an explicit foreach loop.
@@ III-A Code Transformation Knowledge-base
Data Collection We first collect all repositories with $\geq$5 stars on GitHub, whose primary language is C# and had a commit within the last 5 years. We crawl the commit history of the main branch in each of these repos. A commit typically contains a commit message and a changeset or diff representing the difference between current and previous version of files modified within that commit. From these repos, we collect the set of commits whose titles begin with "PERF:" or "[PERF]". This is a convention [19] that’s often used by developers to indicate that their commit contains primarily performance-related changes. This results in $\sim$1500 commits with clean code changes pertaining to performance.
@@ III-B Generating the Prompt
We are given buggy method, $m_{b}$, containing an expensive line of code, $l$, which causes the performance bug. We first abstract the expensive line using Algorithm 1 and get the abstracted pattern $p_{b}$. We then perform a lookup in our knowledge-base using the $p_{b}$, as the query. This results in a list of $K_{p}$ code change and instruction triples, i.e. $[(m_{k},m^{\prime}_{k},t_{k})]_{k=1}^{K_{p}}$, where $m_{k}$ and $m^{\prime}_{k}$ are the before and after versions of a method, where a similar buggy line was fixed and $t_{k}$ is the code transformation instruction constructed based on changes made to the identifiers in the before-after pair. Since we want to find the entry that’s closest to $m_{b}$, we use a code-search technique  [20] to rank each before method, $m_{i}$, based on similarity to $m_{b}$. We take the entry corresponding to the most similar before method, $(m_{j},m^{\prime}_{j},t_{j})$ and construct the prompt using the retrieved code transformation instruction $t_{j}$.
@@ III-C Generating Fixes
Generating the fix itself is relatively simple. We use the constructed prompts to generate completions using the gpt-3.5-turbo [21] model through OpenAI’s REST API. Since it is difficult to tell the model when to stop, it will likely go on outputting code even after it has generated the fixed method. We parse out the completion corresponding to the method by extracting the code until curly braces are balanced and discard the rest of the output.
@@ IV-A Experimental Setup
Dataset We leverage the dataset containing past performance fixes from the DeepDev-PERF study [12]. This dataset consists of 132 instances of performance fixes, which were manually reviewed and confirmed to be performance-related changes by performance experts after examining a broad set of C# commits on GitHub. Furthermore, it was shown that this dataset covers a wide-range of performance bugs. Therefore, a good accuracy over this dataset would demonstrate the effectiveness of a model for this task.Each example in this dataset is of the form $(m,m^{\prime})$, where $m$ and $m^{\prime}$ are the before and after versions, respectively. These are also mostly single method performance improvement fixes, making them ideal for our scenario.
Adding Bug Localization Information Our approach assumes that bug localization information, i.e. the expensive line, is included as part of the input. In practice, one can leverage performance bug localization using profiling data to find the expensive line [22]. However, for the purpose of evaluating our approach, we used a heuristic to localize bugs in the above dataset. We examine the diff between the before and after methods, $m$ and $m^{\prime}$ and take the first line $l_{b}\in m$ that does not match the corresponding line in $m^{\prime}$, when compared in order, as the buggy line.
@@
  • CodeBLEU: Our final metric is CodeBLEU [23], which is a variation of the popular NLP metric, BLEU [24]. In addition to n-gram matching, it also takes code characteristics like abstract syntax tree (AST) and data-flow similarity into account, when comparing programs. Similar to the DeepDev-PERF study, we use the hyper-parameters that had the highest correlation to human scores in the CodeBLEU study, i.e. $\alpha,\beta,\gamma,\delta=0.1,0.1,0.4,0.4$.
@@
  • Closest Match Top-K Accuracy % Due to there being too many suggestions to go through one by one, we use the state of the art code search technique, Aroma [20], to find the suggestion closest to the ground truth. The most similar suggestion is shown to two performance experts, who are then asked whether they consider the model suggestion to be equivalent or better that the developer fix in terms of performance. We consider a suggestion to be correct, if it is found to be correct by both experts. This allows us to get a lower-bound on the model’s actual Top-K accuracy, if one were to exhaustively go through every suggestion.
@@
  • Reasoning-based Prompt Message. Recent work on prompt engineering suggests letting LLMs generate an intermediate reasoning about a problem can greatly improve performance on a given task [16]. As our final variant, we designed a reasoning-based prompt that follows the structure in Figure 5(c). This prompt also includes the buggy method, $m_{b}$, followed by an incomplete comment telling the model what call is being expensive, and coaxing the model to reason about the problem and then generate a possible fix. The main idea is that we first get the model to reason about the problem and come up with a description of the possible fix by completing the prompt message itself, before it generates the fixed code. We do this by giving the model the buggy method, followed by the prompt message. We let the model complete the rest of the comment (i.e. continuing in place of the ellipsis) and then have it generate a fix as well. The prompt follows structure in Figure 5(c).
@@ IV-B Comparing with Other Prompt Variants
Figure 10: An email sent as part of our "In-the-wild" Evaluation. The email contains a fix for a performance bottleneck found using a commercial performance bug detection service in an application. Along with the fix, we also include information regarding the bottleneck, i.e. the bottleneck method and the calling parent method (both highlighted in yellow) and current CPU usage of the bottleneck. At the end of the email, we ask the users to answer a few questions regarding the fix (Section V-A2).
Effect of extracting instruction from retrieved example. Prior work has shown that providing an example in a few shot setting usually results in better performance in a downstream task than providing an instruction[25]. It likely has less to do with the example but more with specificity of the instructions The Retrieval-based One-Shot prompt gives the model the same retrieved before-after pair that the instruction is derived from. However, comparing the results of RAPGen with the One-Shot prompt, we can see that going the extra step of extracting an instruction from the retrieved before-after pair improves the results significantly. We believe that this may be because, given the raw before-after method pair, the model has to perform the diff implicitly and infer the instruction by itself, before applying the change to the unseen method. It is possible for it misunderstand the change and make mistakes in understanding the nature of the fix. We see that our prompts find more exact matches (by $>$10%) and achieves a higher Top-K accuracy score (by $>$20%), compared to retrieval prompts. For the example shown in Figure 9, both the One Shot and RAPGen prompts were able to generate the right fix in our evaluation. However, the One Shot prompt took almost twice as many attempts before arriving at the right fix, compared to RAPGen. The likely reason RAPGen is able to generate a correct suggestion more quickly may be due to a possible fix being readily available as part of the instruction, without the model having to perform in-context learning from a raw code change example and infer the underlying transform.
@@ IV-C Comparison with State-of-the-art
Finally, we compare our approach with the current state of the art approach for performance bug fixing for C#: DeepDev-PERF  [12]. DeepDev-PERF uses a sequence to sequence transformer model, BART [26] and finetunes it for the task of generating performance bugs. The model can output entire patches, as opposed to just completions, allowing it to generate fixes spanning the entire class. In contrast, our approach is intended to fix single method bugs only. Therefore, we only compare our model with DeepDev-PERF’s ability to fix single method fixes, using the dataset they provide.
@@ V-A1 Finding real world problems to fix
We start with 50 codebases, internal to a large company, that are associated with existing services running in the cloud. Each of these services is being profiled by a profiler. The collected profiles are then analyzed by a commercial performance bug detection service that continuously analyzes profiler traces from a given application in the cloud and finds performance bottlenecks.
Each bottleneck is of the form $(b,p)$, where $b$ in the bottleneck method that’s usually an API method and $p$ the calling method from the user’s codebase. We look at the insights generated by the performance bug detection service over CPU profiler traces. For each application, we find bottlenecks $(b,p)$ that meet the following criteria: